# Computing Wasserstein Barycenters through Gradient Flows

### v1

@@ 1 Introduction
Wasserstein barycenters [1] constitute a building block in the analysis of probability measures under Optimal Transport (OT) [2]. This mathematical object extends the notion of the barycenter of points to Wasserstein spaces, thus defining a notion of the average of probability measures. Since it relies on the Wasserstein metric, these barycenters capture the geometry of the underlying space over which the probability measures are defined [3]. As such, they contributed to several areas of machine learning, including model averaging [4], ensembling [5], data augmentation [6], distillation [7, 8], dictionary learning [9], domain adaptation [10, 11, 12], and Bayesian learning [13].
@@
Existing methods for computing Wasserstein barycenters can be divided into three categories. The first category relies on discretization of the input measures and the barycentric measure through empirical measures. These algorithms may be subdivided into fixed-support [14, Algorithm 1][15, 16], and free-support [14, Algorithm 2][11].
@@
The second category relies on parametric models for the input measures, such as Gaussian, or Gaussian mixture models. In the Gaussian case, these barycenters are known as Bures-Wasserstein barycenters [17, 18]. For Gaussian mixtures, [19] shows connections with multi-marginal OT [20] between the components of the input mixtures, while [12] shows that, for components with diagonal covariances, there is an iterative algorithm similar to [14, Algorithm 2].
@@
The third category of methods for computing Wasserstein barycenters relies on neural nets and comes from the effort to scale OT to measures in high-dimensional spaces with a large number of samples. In this sense, [21] proposes parameterizing the barycentric measure through a specific kind of neural net architecture, known as Input Convex Neural Networks (ICNNs) [22]. Furthermore, [23] proposes a bi-level adversarial learning approach that can be used for general ground costs, which is made more robust by semi-unbalanced OT [24]. More recently, [25] proposes computing barycenters with normalizing flows.
@@
Current Wasserstein barycenter algorithms mostly focus on scalability with respect to the number of samples $n$, and dimensions $d$. In this sense, neural networks have a clear advantage over discrete methods, since they can operate at the level of mini-batches. There is another level of scalability, oftentimes overlooked, which is with respect the number of input measures $K$. So far, only [25] has considered experiments in this context. In this regard, methods such as [21, 26] are difficult to scale, as they require $\mathcal{O}(K)$ neural networks to compute the Wasserstein barycenter.
@@
In this paper we propose a new perspective on discrete Wasserstein barycenters, through the lens of Wasserstein Gradient Flows (WGFs) [27, 28], which is an appealing framework for optimization in the space of probability measures. This new understanding of Wasserstein barycenters has several advantages. First, we can scale the discrete barycenter problem, by drawing mini-batches from the input measures. Second, we can scale with respect to the number of input measures, as the representation of the barycentric measure is independent of $K$. Third, based on the established literature on gradient flows, we can define regularized Wasserstein barycenters with suitable functionals on the space of probability measures, which were not possible with previous methods. Lastly, we can modify the underlying metric to integrate label information between samples, leading to barycenters that respect class structure and provide more accurate and semantically meaningful estimates. We show an overview of our approach in Figure 1. Our contributions are:
@@
  • We generalize algorithms from the state-of-the-art [14, 11, 12] under a common framework.
@@
  • We offer a new theoretical result for the mixture-Wasserstein distance between labeled Gaussian mixtures [19, 12].
  • We conduct extensive experiments benchmarking existing methods, including neural network-based Wasserstein barycenter algorithms [21, 23, 24, 25], on both toy datasets and domain adaptation  [^1] In contrast with discrete methods, most experiments in neural network solvers involve generative modeling. We thus establish a bridge between these methods and domain adaptation benchmarks, and show that our approach establishes a new state-of-the-art.
@@ 2.1 Empirical Optimal Transport
OT [2] is a field of mathematics concerned with the transportation of mass at least effort. We refer readers to [29] for a computational treatment of the subject, and [30] for applications in machine learning. Let $P,Q\in\mathcal{P}(\Omega)$, and $c:\Omega^{2}\rightarrow\mathbb{R}$ be a ground-cost. OT was originally founded by Monge [31]. In this formulation, we seek for a mapping $T:\Omega\rightarrow\Omega$ such that,
@@
Alternatively, Kantorovich [32] proposed a formulation in terms of a transport plan $\gamma\in\Gamma(P,Q)$, where $\Gamma(P,Q)$ is the set of all joint measures with marginals $P$ and $Q$. In this case,
@@ 2.2 Gaussian Mixture Optimal Transport
where $\theta=\{(\pi_{i}^{(P)},\mu_{i}^{(P)},\Sigma_{i}^{(P)})\}$ are the parameters of the GMM. The general theory of OT between GMMs was first presented in [19]. The main advantage is that, when the OT plan is restricted to the set of GMMs, i.e, $\gamma\in\Gamma(P,Q)\cap\text{GMM}_{\infty}(d)$, there is a discrete equivalent formulation in terms of the GMM components,
@@
$\omega^{\star}\in\mathbb{R}^{n\times m}$ is a component-to-component OT plan. The term $\mathbb{W}_{2}(P_{i},Q_{j})$ is the Wasserstein distance between Gaussian components, which has closed-form [33] in terms of their parameters,
@@ 2.3 Wasserstein Barycenter and its Variants
In the metric setting, the barycenter problem is known as Fréchet [34] or Karcher [35] means. In our case, given a finite family of measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$ and a set of barycentric coordinates $\lambda\in\Delta_{K}$, we define the Wasserstein barycenters over $(\mathcal{P}_{p,ac}(\Omega),\mathbb{W}_{p})$ through the following optimization problem,
@@
In general, equation 7 does not have a closed-form solution. However, there are algorithms for computing barycenters for $p=2$, when the measures in $\mathcal{P}$ are either empirical measures [14], Gaussian measures [1, 36] or Gaussian mixtures [19, 12], and a further link can be made to multi-marginal OT [20].
The work of [14] is of particular interest to us, since theirs was the first algorithm, based on gradient descent, to optimize equation 7 on empirical measures. The idea is to initialize $z_{0,1}^{(P)},\cdots,z_{0,n}^{(P)}$ randomly (e.g., from a Gaussian measure), and iterate
@@
where $T_{\gamma_{k}}(z_{i}^{(P)})=n\sum_{j=1}^{n}\gamma_{k,i,j}z_{j}^{(Q_{k})}$ is the barycentric map between $Q_{k}$ and $P$. This strategy is reminiscent of the fixed-point approach of [36]. Furthermore, [11] shows that, for feature-label joint measures, the labels can be propagated through [37],
@@ 2.4 Gradient Flows
where $P_{t}$ is a curve in $\mathcal{P}_{p,\text{ac}}(\Omega)$, and $\mathbb{F}:\mathcal{P}_{p,\text{ac}}(\Omega)\rightarrow\mathbb{R}$ is a functional over probability measures. Usually, these functionals take the following form [38],
@@ 3.3 Flows over Joint Measures
In [14], the authors considered barycenters for $p=2$, $\Omega=\mathbb{R}^{d}$, and $d(z,z^{\prime})=\lVert z-z^{\prime}\rVert_{2}$. Therefore, we generalize this setting, since we only need the differentiability of $d$ for our algorithm to work. As we show in our experiments, in machine learning applications one encounters $\Omega=\mathcal{X}\times\mathcal{Y}$, i.e., measures over the joint space of features and labels. Whenever we deal with labeled measures (i.e., $z=(x,y)$), we adopt,
@@
For regression applications (i.e., $\mathcal{Y}=\mathbb{R}$), this distance is quite natural. However, for classification, $\mathcal{Y}$ is categorical (e.g., $\mathcal{Y}=\{1,\cdots,n_{\text{classes}}\}$). One possible strategy, used in [10] and [39], is fixing the labels and flowing only the features. In contrast to these works, we embed $\mathcal{Y}$ into the compact continuous space $\Delta_{K}$, through a one-hot encoding operation. For our flow, we parametrize labels through a change of variables,
@@
where $P_{i,x}=\mathcal{N}(\cdot|\mu_{i}^{(P)},\Sigma_{i}^{(P)})$ is the Gaussian feature marginal. This approach was first proposed in [12], but the authors relied on heuristic arguments to justify this modeling choice. We present the following proposition justifying the label term,
@@
Functionals for joint measures. One of the advantages of our proposed method is regularizing the barycenter calculation with internal, interaction, and potential energy functionals. This idea was already used in practice by [39] for transfer learning problems. Here we propose the following functionals,
@@ 3.4 Convergence
One of the difficulties in analyzing the gradient flow of the functional in equation 14 comes from the fact that $P\mapsto\mathbb{W}_{p}(P,Q)^{p}$ is not geodesically convex in $\mathcal{P}_{p,ac}(\Omega)$ [28, Section 4.4]. As a result, we are minimizing a non-convex functional. Henceforth, we focus on $p=2$, as this is the setting where most results are available. Explicit constants and proofs are available in Section 4 in the Appendix.
Our analysis relies on a measure-theoretic version of the Polyak Łojasiewicz (PL) inequality [40, 41],
@@
which is covered in [42]. Under this assumption, we have the following convergence result,
@@
Let $\Omega=\mathcal{B}(0,R)$ be the closed ball on $\mathbb{R}^{d}$ with radius $R>0$. Let $P^{\star}\in\mathcal{P}_{2,\text{ac}}(\mathcal{B}(0,R))$ be the barycenter of $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K},Q_{k}\in\mathcal{P}_{2,\text{ac}}(\mathcal{B}(0,R))$ $k=1,\cdots,K$, with barycentric coordinates $\lambda=(\lambda_{1},\cdots,\lambda_{K})\in\Delta_{K}$. Assume the inequality 22. Then, the following holds,
@@
Next, we derive a new result for the convergence of the GMM gradient flow (cf. Algorithm 2). Here, we assume $P$ and $Q_{k}\in\mathcal{Q}$ are GMMs. Furthermore, we denote by $\hat{Q}_{k}$ the GMM fitted on data from $Q_{k}$ with the expectation-maximization algorithm. Our result relies on [43, Theorem 3.1]. Therefore, we assume,
@@ 4.1 Toy Example
In Figure 2, we show the Swiss roll measure $Q_{0}$, alongside four variations obtained via $T_{k,\sharp}Q_{0}$, $k=1,\cdots,4$. This example has been used in several state-of-the-art Wasserstein barycenter works [21, 25]. Originally there are labels associated with the samples of the Swiss-roll measure, corresponding to the position of the points in the underlying manifold. We show a summary of our results in Figure 3.
@@
First, we compare unsupervised barycenter solvers in Figures 3 (a) through (h). Quantitatively, empirical methods, notably [14, Algorithm 2] and our WGF algorithm (cf. Algorithm 1) achieve the lowest Wasserstein distance to the ground truth in Figure 2. We note that, overall, while the neural network solvers are usually more scalable than discrete methods in terms of number of samples, their optimization problem is more complicated and very sensitive to hyper-parameters.
(a) CW2B [21].
(b) NOT [23].
(c) Discrete [14].
@@
(g) U-NOT [24].
(h) NormFlow [25].
(i) Discrete [11].
@@ 4.2 Multi-Source Domain Adaptation
Problem Formulation. One of the main applications of Wasserstein barycenters is multi-source domain adaptation (MSDA) [44, 45, 46]. In this setting one needs to adapt multiple labeled source measures $Q_{1},\cdots,Q_{K}$ to a single unlabeled target measure $Q_{T}$. The goal is to learn, from samples $\{\{x_{i}^{(Q_{k})},y_{i}^{(Q_{k})}\}_{i=1}^{n_{k}}\}_{k=1}^{K}$ and $\{x_{i}^{(Q_{T})}\}_{i=1}^{n_{T}}$, a classifier $h$ that achieves low risk in the target domain measure,
@@
for a loss function $\mathcal{L}$ (e.g., cross-entropy loss). We isolate the quality of barycenters by doing adaptation at the level of embeddings. This approach allows us to perform domain adaptation in a higher semantic space, where distributions are more meaningful and comparable across domains. Thus, we assume that a meaningful feature extractor $\phi$, called the backbone, has been previously learned. We obtain the feature extractor by fine-tuning a neural network on the labeled source domain data. In addition, similar to [10, 12], we align the barycenter with the target through OT [47]. More details are available in Section 6 in the Appendix.
@@
Office31 ResNet50 [48] 3287 3 2048 31
BCI-CIV-2a CBraMod [49] 5184 10 200 4
TEP CNN [50] 17289 6 128 29
Office Home ResNet101 [48] 15500 4 2048 65
ISRUC CBraMod [49] 89240 100 512 5
@@
Experimental Setting. We run our experiments on five benchmarks: Office 31 [51], Office Home [52], BCI-CIV-2a [53], ISRUC [54], and TEP [50]. The first two, second two, and last benchmarks correspond to computer vision, neuroscience, and chemical engineering benchmarks, respectively. We show in Table 1 an overview of our experimental setting.
Compared Methods. Overall, we compare seven Wasserstein barycenter strategies with ours. These methods are: discrete barycenters [14, Algorithm 2], normalizing flows [25], continuous 2-Wasserstein barycenters [21], neural OT barycenters [23], and unbalanced neural OT barycenters [24]. We also include labeled barycenter strategies [11, 55]. A brief review of these methods is available in Section 2 in the Appendix. For completeness, we include four other state-of-the-art methods in domain adaptation over embedding vectors. Those methods are: WJDOT [56], DaDiL-R and E [11], and GMM-DaDiL [12]. For each benchmark, we use one domain as the target domain (e.g., Amazon vs. $\{\text{dSLR, Webcam}\}$ in the Office 31 benchamrk), and we measure the classification accuracy, i.e., the percentage of correct predictions.
@@
WJDOT [56] - 86.80 76.59 N/A 76.95 86.13
DaDiL-R [11] - 89.91 77.86 53.41 74.68 86.14
DaDiL-E [11] - 89.79 78.14 N/A 75.89 85.87
GMM-DaDiL [12] - 90.63 78.81 57.10 75.47 86.85
Discrete [14] 81.94 70.42 57.38 70.67 83.81
NormFlow [25] 85.91 76.73 55.45 78.04 82.89
CW2B [21] 86.37 76.44 57.25 75.84 85.83
NOT [23] 86.22 75.36 57.29 76.39 84.93
U-NOT [24] 86.97 76.85 57.17 77.06 85.43
Discrete [11] 87.93 77.09 57.43 78.20 86.09
GMM [12] 88.54 77.87 57.04 74.58 84.67
@@
Main results. We present our main results in Table 2, which reports the average performance per domain on each benchmark. We provide fine-grained results in Section 6 in the Appendix. Overall, labeled barycenter methods (the four last rows in Table 2) have a clear advantage with respect to the unsupervised methods. We argue that label information is pivotal in domain adaptation success, which is consistent with previous research [47, 10, 11, 12]. Among the labeled barycenter methods, our gradient flow framework achieves either the best performance, or remains competitive, surpassing previous methods in MSDA, such as WJDOT and DaDiL in the ISRUC benchmark.
Figure 4: t-SNE [57] visualization of the barycenter of $\text{Art},\text{Product},\text{Real-World}]$ source domains in the Office home benchmark. Colors represent different classes from 11 to $65$. Overall, the classes tend to be more separated when using the repulsion functional $\mathbb{U}$.
Visualization. We now visualize the effect of adding the repulsion interaction energy functional on the Office-Home benchmark. Adding $\mathbb{U}$ leads to classes that are well separated. For computer vision experiments we use ResNets [48] as the backbone, which produce $2048$-dimensional embeddings. For that reason, we use the cosine distance $d(x,x^{\prime})=1-\text{cossim}(x,x^{\prime})$ in equation 21 and we fix $\text{margin}=1$. This choice encourages class clusters to live in orthogonal sub-spaces of the latent space. We show a comparison of the obtained barycenters in Figure 4. Overall, the classes tend to have less overlap when the interaction energy is added to the barycenter functional.
@@
Discrete [14] 70.42 N/A N/A N/A
Discrete [11] 77.09 N/A N/A N/A

### v2

@@ 1 Introduction
Wasserstein barycenters [agueh2011barycenters] constitute a building block in the analysis of probability measures under Optimal Transport (OT) [villani2008optimal]. This mathematical object extends the notion of the barycenter of points to Wasserstein spaces, thus defining a notion of the average of probability measures. Since it relies on the Wasserstein metric, these barycenters capture the geometry of the underlying space over which the probability measures are defined [kloeckner2010geometric]. As such, they contributed to several areas of machine learning, including model averaging [singh2020model], ensembling [coz2023barycenter], data augmentation [zhu2023interpolation], distillation [liu2023dataset, montesuma2024multisource], dictionary learning [schmitz2018wasserstein], domain adaptation [montesuma2021wasserstein, montesuma2023multi, montesuma2024lighter], and Bayesian learning [srivastava2018scalable].
@@
Existing methods for computing Wasserstein barycenters can be divided into three categories. The first category relies on discretization of the input measures and the barycentric measure through empirical measures. These algorithms may be subdivided into fixed-support [cuturi2014fast, Algorithm 1][benamou2015iterative, janati2020debiased], and free-support [cuturi2014fast, Algorithm 2][montesuma2023multi].
@@
The second category relies on parametric models for the input measures, such as Gaussian, or Gaussian mixture models. In the Gaussian case, these barycenters are known as Bures-Wasserstein barycenters [chewi2020gradient, kroshnin2021statistical]. For Gaussian mixtures, [delon2020wasserstein] shows connections with multi-marginal OT [pass2015multi] between the components of the input mixtures, while [montesuma2024lighter] shows that, for components with diagonal covariances, there is an iterative algorithm similar to [cuturi2014fast, Algorithm 2].
@@
The third category of methods for computing Wasserstein barycenters relies on neural nets and comes from the effort to scale OT to measures in high-dimensional spaces with a large number of samples. In this sense, [korotin2021continuous] proposes parameterizing the barycentric measure through a specific kind of neural net architecture, known as Input Convex Neural Networks (ICNNs) [amos2017input]. Furthermore, [kolesov2024estimating] proposes a bi-level adversarial learning approach that can be used for general ground costs, which is made more robust by semi-unbalanced OT [gazdieva2024robust]. More recently, [visentin2025computing] proposes computing barycenters with normalizing flows.
@@
Current Wasserstein barycenter algorithms mostly focus on scalability with respect to the number of samples $n$, and dimensions $d$. In this sense, neural networks have a clear advantage over discrete methods, since they can operate at the level of mini-batches. There is another level of scalability, oftentimes overlooked, which is with respect the number of input measures $K$. So far, only [visentin2025computing] has considered experiments in this context. In this regard, methods such as [korotin2021continuous, korotin2022wasserstein] are difficult to scale, as they require $\mathcal{O}(K)$ neural networks to compute the Wasserstein barycenter.
@@
In this paper we propose a new perspective on discrete Wasserstein barycenters, through the lens of Wasserstein Gradient Flows (WGFs) [ambrosio2008gradient, santambrogio2017euclidean], which is an appealing framework for optimization in the space of probability measures. This new understanding of Wasserstein barycenters has several advantages. First, we can scale the discrete barycenter problem, by drawing mini-batches from the input measures. Second, we can scale with respect to the number of input measures, as the representation of the barycentric measure is independent of $K$. Third, based on the established literature on gradient flows, we can define regularized Wasserstein barycenters with suitable functionals on the space of probability measures, which were not possible with previous methods. Lastly, we can modify the underlying metric to integrate label information between samples, leading to barycenters that respect class structure and provide more accurate and semantically meaningful estimates. We show an overview of our approach in Figure 1. Our contributions are:
@@
  • We generalize algorithms from the state-of-the-art [cuturi2014fast, montesuma2023multi, montesuma2024lighter] under a common framework.
@@
  • We offer a new theoretical result for the mixture-Wasserstein distance between labeled Gaussian mixtures [delon2020wasserstein, montesuma2024lighter].
  • We conduct extensive experiments benchmarking existing methods, including neural network-based Wasserstein barycenter algorithms [korotin2021continuous, kolesov2024estimating, gazdieva2024robust, visentin2025computing], on both toy datasets and domain adaptation  [^1] In contrast with discrete methods, most experiments in neural network solvers involve generative modeling. We thus establish a bridge between these methods and domain adaptation benchmarks, and show that our approach establishes a new state-of-the-art.
@@ 2.1 Empirical Optimal Transport
OT [villani2008optimal] is a field of mathematics concerned with the transportation of mass at least effort. We refer readers to [peyre2019computational] for a computational treatment of the subject, and [montesuma2024recent] for applications in machine learning. Let $P,Q\in\mathcal{P}(\Omega)$, and $c:\Omega^{2}\rightarrow\mathbb{R}$ be a ground-cost. OT was originally founded by Monge [monge1781memoire]. In this formulation, we seek for a mapping $T:\Omega\rightarrow\Omega$ such that,
@@
Alternatively, Kantorovich [kantorovich1942transfer] proposed a formulation in terms of a transport plan $\gamma\in\Gamma(P,Q)$, where $\Gamma(P,Q)$ is the set of all joint measures with marginals $P$ and $Q$. In this case,
@@ 2.2 Gaussian Mixture Optimal Transport
where $\theta=\{(\pi_{i}^{(P)},\mu_{i}^{(P)},\Sigma_{i}^{(P)})\}$ are the parameters of the GMM. The general theory of OT between GMMs was first presented in [delon2020wasserstein]. The main advantage is that, when the OT plan is restricted to the set of GMMs, i.e, $\gamma\in\Gamma(P,Q)\cap\text{GMM}_{\infty}(d)$, there is a discrete equivalent formulation in terms of the GMM components,
@@
$\omega^{\star}\in\mathbb{R}^{n\times m}$ is a component-to-component OT plan. The term $\mathbb{W}_{2}(P_{i},Q_{j})$ is the Wasserstein distance between Gaussian components, which has closed-form [takatsu2011wasserstein] in terms of their parameters,
@@ 2.3 Wasserstein Barycenter and its Variants
In the metric setting, the barycenter problem is known as Fréchet [frechet1948elements] or Karcher [karcher1977riemannian] means. In our case, given a finite family of measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$ and a set of barycentric coordinates $\lambda\in\Delta_{K}$, we define the Wasserstein barycenters over $(\mathcal{P}_{p,ac}(\Omega),\mathbb{W}_{p})$ through the following optimization problem,
@@
In general, equation 7 does not have a closed-form solution. However, there are algorithms for computing barycenters for $p=2$, when the measures in $\mathcal{P}$ are either empirical measures [cuturi2014fast], Gaussian measures [agueh2011barycenters, alvarez2016fixed] or Gaussian mixtures [delon2020wasserstein, montesuma2024lighter], and a further link can be made to multi-marginal OT [pass2015multi].
The work of [cuturi2014fast] is of particular interest to us, since theirs was the first algorithm, based on gradient descent, to optimize equation 7 on empirical measures. The idea is to initialize $z_{0,1}^{(P)},\cdots,z_{0,n}^{(P)}$ randomly (e.g., from a Gaussian measure), and iterate
@@
where $T_{\gamma_{k}}(z_{i}^{(P)})=n\sum_{j=1}^{n}\gamma_{k,i,j}z_{j}^{(Q_{k})}$ is the barycentric map between $Q_{k}$ and $P$. This strategy is reminiscent of the fixed-point approach of [alvarez2016fixed]. Furthermore, [montesuma2023multi] shows that, for feature-label joint measures, the labels can be propagated through [redko2019optimal],
@@ 2.4 Gradient Flows
where $P_{t}$ is a curve in $\mathcal{P}_{p,\text{ac}}(\Omega)$, and $\mathbb{F}:\mathcal{P}_{p,\text{ac}}(\Omega)\rightarrow\mathbb{R}$ is a functional over probability measures. Usually, these functionals take the following form [santambrogio2015optimal],
@@ 3.3 Flows over Joint Measures
In [cuturi2014fast], the authors considered barycenters for $p=2$, $\Omega=\mathbb{R}^{d}$, and $d(z,z^{\prime})=\lVert z-z^{\prime}\rVert_{2}$. Therefore, we generalize this setting, since we only need the differentiability of $d$ for our algorithm to work. As we show in our experiments, in machine learning applications one encounters $\Omega=\mathcal{X}\times\mathcal{Y}$, i.e., measures over the joint space of features and labels. Whenever we deal with labeled measures (i.e., $z=(x,y)$), we adopt,
@@
For regression applications (i.e., $\mathcal{Y}=\mathbb{R}$), this distance is quite natural. However, for classification, $\mathcal{Y}$ is categorical (e.g., $\mathcal{Y}=\{1,\cdots,n_{\text{classes}}\}$). One possible strategy, used in [montesuma2021wasserstein] and [alvarez2021dataset], is fixing the labels and flowing only the features. In contrast to these works, we embed $\mathcal{Y}$ into the compact continuous space $\Delta_{K}$, through a one-hot encoding operation. For our flow, we parametrize labels through a change of variables,
@@
where $P_{i,x}=\mathcal{N}(\cdot|\mu_{i}^{(P)},\Sigma_{i}^{(P)})$ is the Gaussian feature marginal. This approach was first proposed in [montesuma2024lighter], but the authors relied on heuristic arguments to justify this modeling choice. We present the following proposition justifying the label term,
@@
Functionals for joint measures. One of the advantages of our proposed method is regularizing the barycenter calculation with internal, interaction, and potential energy functionals. This idea was already used in practice by [alvarez2021dataset] for transfer learning problems. Here we propose the following functionals,
@@ 3.4 Convergence
One of the difficulties in analyzing the gradient flow of the functional in equation 14 comes from the fact that $P\mapsto\mathbb{W}_{p}(P,Q)^{p}$ is not geodesically convex in $\mathcal{P}_{p,ac}(\Omega)$ [santambrogio2017euclidean, Section 4.4]. As a result, we are minimizing a non-convex functional. Henceforth, we focus on $p=2$, as this is the setting where most results are available. Explicit constants and proofs are available in Section 4 in the Appendix.
Our analysis relies on a measure-theoretic version of the Polyak Łojasiewicz (PL) inequality [polyak1964gradient, karimi2016linear],
@@
which is covered in [chewi2024statistical]. Under this assumption, we have the following convergence result,
@@
Let $\Omega=\mathcal{B}(0,R)$ be the closed ball on $\mathbb{R}^{d}$ with radius $R>0$. Let $P^{\star}\in\mathcal{P}_{2,\text{ac}}(\mathcal{B}(0,R))$ be the barycenter of $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K},Q_{k}\in\mathcal{P}_{2,\text{ac}}(\mathcal{B}(0,R))$ $k=1,\cdots,K$, with barycentric coordinates $\lambda=(\lambda_{1},\cdots,\lambda_{K})\in\Delta_{K}$. Assume the inequality 47. Then, the following holds,
@@
Next, we derive a new result for the convergence of the GMM gradient flow (cf. Algorithm 2). Here, we assume $P$ and $Q_{k}\in\mathcal{Q}$ are GMMs. Furthermore, we denote by $\hat{Q}_{k}$ the GMM fitted on data from $Q_{k}$ with the expectation-maximization algorithm. Our result relies on [boite2025differentiable, Theorem 3.1]. Therefore, we assume,
@@ 4.1 Toy Example
In Figure 2, we show the Swiss roll measure $Q_{0}$, alongside four variations obtained via $T_{k,\sharp}Q_{0}$, $k=1,\cdots,4$. This example has been used in several state-of-the-art Wasserstein barycenter works [korotin2021continuous, visentin2025computing]. Originally there are labels associated with the samples of the Swiss-roll measure, corresponding to the position of the points in the underlying manifold. We show a summary of our results in Figure 3.
@@
First, we compare unsupervised barycenter solvers in Figures 3 (a) through (h). Quantitatively, empirical methods, notably [cuturi2014fast, Algorithm 2] and our WGF algorithm (cf. Algorithm 1) achieve the lowest Wasserstein distance to the ground truth in Figure 2. We note that, overall, while the neural network solvers are usually more scalable than discrete methods in terms of number of samples, their optimization problem is more complicated and very sensitive to hyper-parameters.
(a) CW2B [korotin2021continuous].
(b) NOT [kolesov2024estimating].
(c) Discrete [cuturi2014fast].
@@
(g) U-NOT [gazdieva2024robust].
(h) NormFlow [visentin2025computing].
(i) Discrete [montesuma2023multi].
@@ 4.2 Multi-Source Domain Adaptation
Problem Formulation. One of the main applications of Wasserstein barycenters is multi-source domain adaptation (MSDA) [kouw2019review, sun2015survey, pan2009survey]. In this setting one needs to adapt multiple labeled source measures $Q_{1},\cdots,Q_{K}$ to a single unlabeled target measure $Q_{T}$. The goal is to learn, from samples $\{\{x_{i}^{(Q_{k})},y_{i}^{(Q_{k})}\}_{i=1}^{n_{k}}\}_{k=1}^{K}$ and $\{x_{i}^{(Q_{T})}\}_{i=1}^{n_{T}}$, a classifier $h$ that achieves low risk in the target domain measure,
@@
for a loss function $\mathcal{L}$ (e.g., cross-entropy loss). We isolate the quality of barycenters by doing adaptation at the level of embeddings. This approach allows us to perform domain adaptation in a higher semantic space, where distributions are more meaningful and comparable across domains. Thus, we assume that a meaningful feature extractor $\phi$, called the backbone, has been previously learned. We obtain the feature extractor by fine-tuning a neural network on the labeled source domain data. In addition, similar to [montesuma2021wasserstein, montesuma2024lighter], we align the barycenter with the target through OT [courty2016optimal]. More details are available in Section 6 in the Appendix.
@@
Office31 ResNet50 [he2016deep] 3287 3 2048 31
BCI-CIV-2a CBraMod [wang2024cbramod] 5184 10 200 4
TEP CNN [montesuma2024benchmarking] 17289 6 128 29
Office Home ResNet101 [he2016deep] 15500 4 2048 65
ISRUC CBraMod [wang2024cbramod] 89240 100 512 5
@@
Experimental Setting. We run our experiments on five benchmarks: Office 31 [saenko2010adapting], Office Home [venkateswara2017deep], BCI-CIV-2a [brunner2008bci], ISRUC [khalighi2013automatic], and TEP [montesuma2024benchmarking]. The first two, second two, and last benchmarks correspond to computer vision, neuroscience, and chemical engineering benchmarks, respectively. We show in Table 1 an overview of our experimental setting.
Compared Methods. Overall, we compare seven Wasserstein barycenter strategies with ours. These methods are: discrete barycenters [cuturi2014fast, Algorithm 2], normalizing flows [visentin2025computing], continuous 2-Wasserstein barycenters [korotin2021continuous], neural OT barycenters [kolesov2024estimating], and unbalanced neural OT barycenters [gazdieva2024robust]. We also include labeled barycenter strategies [montesuma2023multi, montesuma2024multi]. A brief review of these methods is available in Section 2 in the Appendix. For completeness, we include four other state-of-the-art methods in domain adaptation over embedding vectors. Those methods are: WJDOT [turrisi2022multi], DaDiL-R and E [montesuma2023multi], and GMM-DaDiL [montesuma2024lighter]. For each benchmark, we use one domain as the target domain (e.g., Amazon vs. $\{\text{dSLR, Webcam}\}$ in the Office 31 benchamrk), and we measure the classification accuracy, i.e., the percentage of correct predictions.
@@
WJDOT [turrisi2022multi] - 86.80 76.59 N/A 76.95 86.13
DaDiL-R [montesuma2023multi] - 89.91 77.86 53.41 74.68 86.14
DaDiL-E [montesuma2023multi] - 89.79 78.14 N/A 75.89 85.87
GMM-DaDiL [montesuma2024lighter] - 90.63 78.81 57.10 75.47 86.85
Discrete [cuturi2014fast] 81.94 70.42 57.38 70.67 83.81
NormFlow [visentin2025computing] 85.91 76.73 55.45 78.04 82.89
CW2B [korotin2021continuous] 86.37 76.44 57.25 75.84 85.83
NOT [kolesov2024estimating] 86.22 75.36 57.29 76.39 84.93
U-NOT [gazdieva2024robust] 86.97 76.85 57.17 77.06 85.43
Discrete [montesuma2023multi] 87.93 77.09 57.43 78.20 86.09
GMM [montesuma2024lighter] 88.54 77.87 57.04 74.58 84.67
@@
Main results. We present our main results in Table 2, which reports the average performance per domain on each benchmark. We provide fine-grained results in Section 6 in the Appendix. Overall, labeled barycenter methods (the four last rows in Table 2) have a clear advantage with respect to the unsupervised methods. We argue that label information is pivotal in domain adaptation success, which is consistent with previous research [courty2016optimal, montesuma2021wasserstein, montesuma2023multi, montesuma2024lighter]. Among the labeled barycenter methods, our gradient flow framework achieves either the best performance, or remains competitive, surpassing previous methods in MSDA, such as WJDOT and DaDiL in the ISRUC benchmark.
Figure 4: t-SNE [maaten2008visualizing] visualization of the barycenter of $\text{Art},\text{Product},\text{Real-World}]$ source domains in the Office home benchmark. Colors represent different classes from $1$ to $65$. Overall, the classes tend to be more separated when using the repulsion functional $\mathbb{U}$.
Visualization. We now visualize the effect of adding the repulsion interaction energy functional on the Office-Home benchmark. Adding $\mathbb{U}$ leads to classes that are well separated. For computer vision experiments we use ResNets [he2016deep] as the backbone, which produce $2048$-dimensional embeddings. For that reason, we use the cosine distance $d(x,x^{\prime})=1-\text{cossim}(x,x^{\prime})$ in equation 21 and we fix $\text{margin}=1$. This choice encourages class clusters to live in orthogonal sub-spaces of the latent space. We show a comparison of the obtained barycenters in Figure 4. Overall, the classes tend to have less overlap when the interaction energy is added to the barycenter functional.
@@
Discrete [cuturi2014fast] 70.42 N/A N/A N/A
Discrete [montesuma2023multi] 77.09 N/A N/A N/A